# SPIRAL: Self-Play on Zero-Sum Games Incentivizes Reasoning via Multi-Agent Multi-Turn Reinforcement Learning

### v2

@@ 4.4 Implementation
To implement SPIRAL, we develop a truly online multi-agent, multi-turn RL system for finetuning LLMs. Our training framework builds on Oat (Liu et al., 2024), which provides interfaces of a distributed actor-learner architecture (Espeholt et al., 2018). We instantiate actors to execute the self-play loop, using vLLM (Kwon et al., 2023) for efficient model inference and TextArena (Guertler et al., 2025) to simulate the language games. The resulting multi-turn, multi-game self-play experiences are used to update the LLM via policy gradient methods (Sutton and Barto, 2018), incorporating our proposed Role-conditioned Advantage Estimation in the collocated learner.
5 Experiment Setup
We design experiments to carefully test how training on different zero-sum games can improve reasoning capabilities that generalize to novel downstream math and reasoning tasks, as well as novel zero-sum games. This section introduces the selection of game environments (Sec. 5.1) and the training (Sec. 5.2) and evaluation (Sec. 5.3) settings.
We investigate four key questions:
  1. 1.
    RQ1: Can self-play on zero-sum games improve math and general reasoning capabilities? We test whether competitive environment alone can develop transferable reasoning capabilities without domain-specific training data, and analyze emergent patterns to understand transfer mechanisms.
  2. 2.
    RQ2: Does self-play’s automatic curriculum outperform fixed-opponent training? We evaluate whether the automatic curriculum with adaptive difficulty generated through self-play produces more robust reasoning than training against static opponents.
  3. 3.
    RQ3: Do different games develop specialized reasoning skills? We test whether distinct game mechanics cultivate complementary cognitive abilities and whether multi-game training creates synergistic benefits.
  4. 4.
    RQ4: Is Role-Conditioned Advantage Estimation (RAE) essential for stable self-play training? We ablate our variance reduction technique to determine whether RAE is necessary to stabilize training and prevent thinking collapse in zero-sum self-play with shared parameters.
5.1 Game Environments
We select three games from TextArena (Guertler et al., 2025) with distinct cognitive requirements:
TicTacToe (Spatial Reasoning) (Beck, 2008). A 3×3 grid game requiring pattern recognition and adversarial planning. Players must identify winning configurations, block opponent threats, and plan multi-step forcing sequences. We hypothesize these skills transfer to geometric problem-solving and spatial visualization tasks. Tic-Tac-Toe is a deterministic, perfect-information game, which isolates pure strategic reasoning from uncertainty management.
Kuhn Poker (Probabilistic Reasoning) (Kuhn, 1950). A minimal poker variant with three cards (Jack, Queen, King) where players bet with hidden information. Success requires probability calculation, opponent modeling, and decision-making under uncertainty. We expect these capabilities to transfer to problems involving probability, expected value, and strategic uncertainty.
Simple Negotiation (Strategic Optimization) (Nash, 1950). A resource trading game where two players exchange Wood and Gold with opposing valuations to maximize portfolio value. Success requires multi-step planning, theory of mind to model opponent preferences, and strategic communication through proposals and counteroffers. We hypothesize these skills transfer to optimization problems, resource allocation tasks, and multi-constraint reasoning that requires balancing competing objectives. Please see Fig. 7 in App. B for example environment observations.
5.2 Training Settings
We employ Qwen3-4B-Base (Yang et al., 2025) as the base model for all experiments. The training spans 400 steps with 128 samples per step, totaling 51,200 game transitions. We use a sampling temperature $\tau=1.0$ to collect game experiences and Adam (Kingma and Ba, 2014) with a constant learning rate of $1\times 10^{-6}$ to optimize the model. The training batch size is set as 128 and the discount factor $\gamma=1.0$ (no discounting). We use $\alpha=0.95$ for the exponential moving average decay of role-specific baselines. Detailed hyperparameters are listed in Table 7 in App. B. Each experiment takes around 25 hours for full parameter tuning on 8 H100 GPUs with our distributed actor-learner architecture.
5.3 Evaluation Metrics
We track three key metrics throughout the training process: the frequency of invalid moves, the game lengths and the win rates against Gemini-2.0-Flash-Lite [^3] Accessed via https://openrouter.ai/google/gemini-2.0-flash-lite-001., to monitor the learning dynamics and determine the best models. To evaluate the trained models, we test their win rates against a fixed opponent on both training and unseen games and standard reasoning benchmarks, as detailed below.
Out-of-Distribution Generalization to Unseen Games. We evaluate transfer learning on seven unseen games that test whether learned skills generalize beyond training domains: Snake and Connect Four (spatial reasoning), Pig Dice and Liar’s Dice (probabilistic reasoning), Truth and Deception (strategic optimization). These games specifically probe whether spatial reasoning from TicTacToe, probabilistic reasoning from Kuhn Poker, and strategic optimization from Simple Negotiation transfer to novel game mechanics. We report the average win rate of our models against Gemini-2.0-Flash [^4] Accessed via https://openrouter.ai/google/gemini-2.0-flash-001. across 512 independent games, with the starting player randomized in each match to eliminate any first-move advantage.
Transfer to Standard Reasoning Benchmarks. To investigate whether the reasoning abilities developed through gameplay could transfer to non-game contexts, we evaluate our models on a suite of established benchmarks. For mathematical reasoning, we use MATH500 (Hendrycks et al., 2021), OlympiadBench (He et al., 2024), Minerva Math (Lewkowycz et al., 2022), AIME24, AIME25, and AMC23 datasets, which cover a wide range of topics including algebra, geometry, and competitive mathematics. For general reasoning, we utilize GPQA (Rein et al., 2024), which consists of graduate-level science questions, and MMLU-Pro (Wang et al., 2024), a benchmark for multidisciplinary knowledge. All evaluations on these benchmarks are conducted in a zero-shot setting [^5] Except for the base model, for which we provide few-shot examples that follow the Qwen3 Report settings. to determine if game-induced reasoning could be successfully transferred to general problem-solving. We use sampling temperature $0.6$, top-p $0.95$ and $4$ rollouts for all evaluations.
Reasoning Pattern Analysis. We analyze the cognitive patterns that emerge during training and their transfer to mathematical reasoning. Using GPT-4.1 (gpt-4.1-2025-04-14(OpenAI, 2025) as LLM-as-Judge (Zheng et al., 2023), we classify reasoning traces from 290 game trajectories and 46,792 math problem solutions into three core patterns: Case-by-Case Analysis (systematic enumeration of scenarios), Expected Value Calculation (probabilistic decision-making), and Pattern Recognition (identifying regularities and structures). The complete prompts used for pattern discovery, classification, and transfer analysis are provided in App. D. We track the evolution of these patterns across training checkpoints (early: step 0, mid: step 128, late: step 400) in both game and math domains to measure transfer rates. This analysis reveals which game-induced reasoning strategies generalize to academic problem-solving and explains the mechanism behind cross-domain transfer.
6 Experimental Results and Findings
6.1 RQ1: Can self-play on zero-sum games improve math and general reasoning capabilities?
Table 1 presents our main results. SPIRAL, trained exclusively on Kuhn Poker through self-play, achieves notable improvements on both mathematical and general reasoning benchmarks. Notably, it even surpasses a model that was fine-tuned on 25,000 winning game trajectories from Qwen3-32B playing against itself.
Table 1: Reasoning benchmark performance. SPIRAL improves reasoning without any domain-specific training data.
@@ 6.1 RQ1: Can self-play on zero-sum games improve math and general reasoning capabilities?
Model MATH500 AIME’24 AIME’25 OlympiadBench AMC-23 Minerva Math GPQA MMLU-Pro Average
Qwen3-4B-Base 65.8 10.0 3.3 33.3 50.0 24.3 30.6 47.2 32.9
SFT (Qwen3-32B Distill) 76.0 16.7 13.3 38.4 50.0 31.2 33.0 48.8 39.7
SPIRAL (Ours) 76.4 13.3 10.0 38.4 57.5 42.4 37.0 57.7 41.6
Improvement +10.6 +3.3 +6.7 +5.1 +7.5 +18.1 +6.4 +10.5 +8.7
Few-shot evaluation following Qwen3 technical report.
The results show consistent improvements across mathematical reasoning (+10.6% MATH500, +6.7% AIME’25, +7.5% AMC-23), and general reasoning (+6.4% GPQA, +10.5% MMLU-Pro). Notably, SPIRAL achieves these gains despite never seeing a single mathematical equation, academic problem, or domain-specific training example during its entire training process. The 18.1% improvement on Minerva Math and overall 8.7% gain across math benchmarks demonstrate that competitive game environments successfully force discovery of reasoning patterns that transfer broadly across domains.
6.1.1 Case Study: How Game Strategies Transfer to Mathematical Reasoning
To understand this surprising transfer, we use GPT-4.1 (gpt-4.1-2025-04-14) to analyze 290 game trajectories and 46,792 math solutions across training checkpoints, identifying three core reasoning patterns that emerge during gameplay and transfer to math reasoning domains.
Pattern Discovery and Examples. We first illustrate what these patterns look like in both game and mathematical contexts:
@@ 6.1.1 Case Study: How Game Strategies Transfer to Mathematical Reasoning
Table 2: Representative reasoning patterns and their manifestation in games and mathematics. We use an LLM-as-a-judge framework to automatically evaluate the frequency of these patterns over training.
@@
Case-by-Case Analysis <think>
@@
Since losing 1 chip (fold) is better than losing 2 chips (call and lose), the best action is to fold.
</think>
<think> To solve $m^{2}+2\cdot 3^{n}=m(2^{n+1}-1)$:
@@
Case $n\geqslant 4$: No solutions (exponential dominance) </think>
Expected Value Calculation <think> With card J against opponent’s bet:
@@
Since EV(fold) > EV(call), fold is optimal. </think>
<think> For average of $A+BC+DEF$ over permutations:
@@
Total = $-\frac{1}{3}+\frac{1}{9}-\frac{1}{27}=-\frac{7}{27}$ </think>
Pattern Recognition <think> Previous rounds: Player 0 had K both times and won. Pattern identified: Player 0 likely has strong cards or bluffs consistently. Given I have J (weak), betting might exploit their aggressive calling pattern. </think> <think> Sum of three consecutive integers $(n-1)+n+(n+1)=3n$.
@@
Smallest: $k=3\Rightarrow 3n=27$ </think>
Pattern Evolution and Transfer Analysis. Using GPT-4.1 to classify reasoning strategies, we tracked how these patterns develop during training and transfer to mathematics:
Figure 4: Evolution of reasoning patterns during SPIRAL training and their transfer to mathematical reasoning. We track three core reasoning patterns (Pattern Recognition, Expected Value Calculation, and Case-by-Case Analysis) across 290 game trajectories and 46,792 math solutions. Left: In the game domain, all patterns show substantial growth, with Expected Value Calculation reaching 78% by late training. Middle: These patterns transfer to mathematical reasoning with varying effectiveness: Case-by-Case Analysis maintains high transfer (72% to 71%), Pattern Recognition shows amplification (35% to 45%), while Expected Value Calculation transfers more selectively (78% to 28%). Right: Math benchmark scores improve from 31.2 to 39.6 as these reasoning patterns develop, demonstrating that game-learned strategies enhance mathematical problem-solving capabilities.
Fig. 4 shows how reasoning patterns evolve in both game and math domains across training checkpoints. These patterns develop at different rates and transfer with varying effectiveness:
Case-by-Case Analysis: We find that the meta-cognitive skill of systematic enumeration, or breaking problems down into cases, transfers highly effectively from strategic games to math problems (as shown in Fig. 4). We hypothesize this is because it represents a domain-agnostic way of structuring thinking that generally improves reasoning performance. Whether analyzing opponent possibilities in Poker or solution branches in mathematics, the core skill is the same.
Expected Value Calculation: We find that training on games increases the use of expected value calculation from 15% to 78% in the games themselves, and from 12% to 28% in the math domain. While game-specific probabilistic reasoning is more frequent in games than in the math benchmark (likely because most math problems lack explicit decision-theoretic structure), we find that this form of reasoning benefits probability and optimization problems where EV concepts directly apply.
Pattern Recognition: Interestingly, this form of reasoning shows an amplification effect, where the proportion evident in the math domain actually exceeds that of the game domain. We hypothesize that because mathematics inherently requires pattern recognition, game training enhances an already-present mathematical skill which is then deployed even more frequently when the model is prompted to answer a math problem. The combined training produces stronger pattern recognition (45%) than games alone require (35%).
Transfer Mechanism. We believe three factors enable this cross-domain transfer:
1. Competitive pressure strips away memorization: Self-play opponents continuously evolve, forcing models to develop genuine reasoning rather than pattern matching.
2. Games isolate pure reasoning: Without domain complexity, games teach fundamental cognitive operations (enumeration, evaluation, synthesis) that generalize effectively.
3. Structured output bridges domains: The <think> format learned in games provides a reasoning scaffold that models reuse for mathematical problems.
Finding 1: Zero-sum games teach transferable reasoning without domain data Self-play on Kuhn Poker improves math and general reasoning benchmarks despite never seeing benchmark related problems. This works because gameplay forces discovery of three reasoning patterns: Case-by-Case Analysis, Expected Value Calculation, and Pattern Recognition. These skills are developed during competing against an evolving opponent.
6.2 RQ2: Does self-play’s automatic curriculum outperform fixed-opponent training?
Although self-play promises potentially infinite curriculum generation, it remains unclear whether the evolving challenge actually produces better learning than simply training against fixed opponents. We test whether adaptive difficulty through self-play leads to more robust reasoning development.
Experimental Setup. We compare four training settings on the Kuhn Poker environment: (1) Self-Play: the model trains against continuously updating copies of itself; (2) Random Opponent: the opponent always provides valid actions but poses minimal strategic challenge; (3) Mistral Opponent: using Mistral-Small-3 (Mistral, 2025) as a fixed intermediate-level opponent; and (4) Gemini Opponent,: using Gemini-2.0-Flash-Lite (Gemini Team, Google, 2025) as a fixed strong opponent. All variants use identical hyperparameters, infrastructure, and 400 training steps.
Results Analysis.
Figure 5 reveals two distinct failure modes in fixed-opponent training:
Figure 5: Performance comparison of self-play training and fixed-opponent baselines. All evaluations are averaged over multiple games/benchmarks (see Sec. 5.3). Mistral Opponent refers to against Mistral-Small-3; Gemini Opponent refers to against Gemini-2.0-Flash-Lite.
1. Curse of Turns in Format Learning: Training against random opponents leads to complete collapse. Although these rule-based random agents offer no strategic challenge, they always produce valid moves. Consequently, the model must generate correctly formatted, valid actions at every turn throughout the trajectory in order to receive any positive reward. However, the probability of producing a fully valid trajectory decreases exponentially with the episode length, making it extremely difficult to explore and learn the correct action format over long horizons.
2. Exploitation of Static Strategies: Fixed model-based opponents (Mistral, Gemini) facilitate format learning more effectively than random opponents, as they occasionally produce invalid moves that allow the agent to win and receive reward signals. However, we find that training against these static opponents often leads to overfitting to their fixed strategies, resulting in poor generalization to math and general reasoning benchmarks (middle and right of Fig. 5).
Continuous Adaptation Through Self-Play. Self-play uniquely maintains learning momentum throughout training and generalizes better because self-play automatically adjusts difficulty. As shown in Table 3, training against a fixed opponent (Gemini) initially gets zero win rate (thus no positive reward), leading to no effective learning before step 128 (see left of Fig. 5); at step 384 the agent learns to outperform the fixed opponent with $62.5\%$ win rate, gradually exploiting its static strategies. On the other hand, against previous versions of itself, self-play maintains 50-52% win rates throughout training, confirming the opponent evolves to match current capabilities.
@@ 6.2 RQ2: Does self-play’s automatic curriculum outperform fixed-opponent training?
The reasoning transfer results are even more striking. Self-play achieves 40% on math reasoning and 45% on general reasoning, outperforming the best fixed opponent (Gemini) by 5 and 3 percentage points respectively. The relative improvement demonstrates that the diverse strategies required by an evolving opponent create more generalizable reasoning patterns than exploiting static weaknesses.
Finding 2: Self-play creates adaptive curriculum that outperforms fixed opponents Self-play achieves better reasoning transfer than any fixed-opponent training. Rule-based random opponents cause complete collapse due to the curse of turns in format learning. Training against model-based opponents plateaus once the agent finds winning counter-strategies. Self-play avoids both failures by continuously adjusting difficulty: the opponent improves as the agent improves, forcing ongoing adaptation rather than exploitation of static weaknesses.
6.3 RQ3: Do different games develop specialized reasoning skills?
We investigate whether distinct game mechanics cultivate different cognitive abilities. If games truly serve as reasoning gymnasiums, then training on TicTacToe should develop different skills than training on Kuhn Poker or Simple Negotiation. Furthermore, we test whether combining multiple games produces synergistic benefits beyond the sum of individual training.
Experimental Design.
To isolate game-specific skill development, we train separate models exclusively on each game using identical hyperparameters and training duration. We then evaluate these specialists in three ways: (1) performance on their training game, (2) transfer to out-of-distribution games, and (3) generalization to math and general reasoning benchmarks. This design reveals whether games develop narrow tactics or transferable cognitive capabilities. For game performance evaluation, we use two complementary approaches: head-to-head competition between specialists to directly compare learned strategies, and performance against a fixed strong opponent (Gemini-2.0-Flash-Lite) to measure absolute skill levels. This dual evaluation reveals both relative strengths between specialists and their absolute capabilities.
Head-to-Head Competition Reveals Specialized Strategies.
We first examine how specialists perform when competing directly against each other. This head-to-head evaluation uses games that isolate specific cognitive skills: we pair each training game with an unseen game requiring similar abilities (Snake for spatial reasoning like TicTacToe, Pig Dice for probabilistic reasoning like Kuhn Poker, and Truth and Deception for strategic optimization like Simple Negotiation).
@@ 6.3 RQ3: Do different games develop specialized reasoning skills?
Models trained exclusively on TicTacToe excel at spatial pattern recognition, achieving 57.5% win rate against other specialists on their training game. When tested on Snake, an unseen spatial reasoning game, they maintain 56.0% performance, demonstrating robust transfer. Kuhn Poker specialists develop probabilistic reasoning, dominating their training game (64.2%) and remarkably achieving 91.7% on Pig Dice. Simple Negotiation training produces models skilled in strategic optimization (62.7% on training game, 55.8% on Truth and Deception).
Multi-Game Training Creates Synergistic Benefits.
While head-to-head competition reveals specialized strategies, we also evaluate absolute performance against a strong fixed opponent (Gemini-2.0-Flash-Lite) across both training games and randomly sampled out-of-distribution games. The multi-game model demonstrates broader capabilities that often exceed specialist performance:
Table 5: Multi-game training achieves competitive performance across all training games while excelling at novel composite challenges. All win rates shown are against Gemini-2.0-Flash-Lite as a fixed opponent. The multi-game model outperforms all specialists on average, demonstrating that diverse game training develops more flexible reasoning.
@@
Model TicTacToe KuhnPoker Simple Negotiation Liar’s Dice Connect Four Snake Average
Base Model 17.5% 21.5% 15.6% 16.7% 20.5% 20.6% 18.7%

### v3

@@ 3 The SPIRAL Framework
Implementation. To implement SPIRAL, we develop a truly online multi-agent, multi-turn RL system for finetuning LLMs. Our training framework builds on Oat (Liu et al., 2024), which provides interfaces of a distributed actor-learner architecture (Espeholt et al., 2018). We instantiate actors to execute the self-play loop, using vLLM (Kwon et al., 2023) for efficient model inference and TextArena (Guertler et al., 2025) to simulate the language games. The resulting multi-turn, multi-game self-play experiences are used to update the LLM via policy gradient methods (Sutton & Barto, 2018), incorporating our proposed Role-conditioned Advantage Estimation in the collocated learner.
4 Experimental Results
Table 1: Reasoning benchmark performance. The “-Kuhn” suffix denotes fine-tuning solely on a single game (Kuhn Poker), while the “-Multi” suffix indicates fine-tuning on all three games. SPIRAL improves reasoning without any domain-specific training data. Few-shot evaluation following Qwen3 technical report.
@@ 4 Experimental Results
Model Math500 AIME24 AIME25 Olympiad AMC-23 Minerva GPQA-D MMLU-Pro Average
Qwen3-4B-Base $73.4$ $9.6$ $6.2$ $33.3$ $42.4$ $29.4$ $30.6$ $47.2$ $34.0$
   + SFT-Kuhn $74.0$ $11.0$ $10.4$ $36.7$ $48.6$ $36.8$ $33.0$ $48.8$ $37.4$
   + SFT-Multi $74.2$ $13.7$ $11.7$ $37.6$ $51.1$ $40.1$ $37.8$ $51.3$ $39.7$
   + SPIRAL-Kuhn (Ours) $76.4$ $18.2$ $15.6$ $38.4$ $61.2$ $42.4$ $37.0$ $57.7$ $43.4$
   + SPIRAL-Multi (Ours) 78.278.2+4.8 19.719.7+10.1 13.313.3+7.1 41.841.8+8.5 61.661.6+19.2 42.642.6+13.2 40.140.1+9.5 58.558.5+11.3 44.544.5+10.5
Qwen3-8B-Base $77.0$ $12.1$ $11.2$ $33.5$ $50.6$ $38.2$ $38.0$ $55.7$ $39.5$
   + SFT-Multi $82.8$ $19.9$ $15.6$ $45.9$ $63.5$ $40.8$ $41.6$ $58.8$ $46.1$
   + SPIRAL-Multi (Ours) 86.686.6+9.6 26.226.2+14.1 16.816.8+5.6 49.649.6+16.1 65.265.2+14.6 46.346.3+8.1 44.644.6+6.6 61.161.1+5.4 49.649.6+10.1
Octothinker-8B-Base $65.6$ $1.7$ $0.5$ $26.6$ $33.5$ $25.7$ $22.1$ $30.8$ $25.8$
   + SFT-Multi $66.0$ $3.3$ $3.8$ $23.9$ $31.0$ $23.8$ $24.9$ $39.1$ $27.0$
   + SPIRAL-Multi (Ours) 68.668.6+3.0 5.35.3+3.6 4.84.8+4.3 33.733.7+7.1 43.243.2+9.7 32.032.0+6.3 33.833.8+11.7 49.349.3+18.5 33.833.8+8.0
Llama-3.1-8B-Instruct $46.4$ $4.6$ $0.7$ $13.8$ $23.3$ $22.8$ $30.2$ $49.1$ $23.9$
   + SFT-Multi $51.8$ $4.6$ $0.7$ 19.1 $23.3$ $21.7$ $30.0$ $48.9$ $25.0$
   + SPIRAL-Multi (Ours) 49.849.8+3.4 4.94.9+0.3 1.81.8+1.1 17.317.3+3.5 26.026.0+2.7 24.624.6+1.8 32.232.2+2.0 50.450.4+1.3 25.925.9+2.0
DeepSeek-Distill-Qwen-7B $90.8$ $53.0$ $39.5$ $56.9$ $89.3$ $48.2$ $48.6$ $57.1$ $60.4$
   + SFT-Multi $91.8$ $49.3$ $36.6$ $52.4$ $88.2$ $48.2$ $44.5$ $55.6$ $58.3$
   + SPIRAL-Multi (Ours) 93.093.0+2.2 54.154.1+1.1 40.840.8+1.3 57.957.9+1.0 89.389.3+0.0 51.151.1+2.9 49.649.6+1.0 58.958.9+1.8 61.861.8+1.4
We evaluate SPIRAL across diverse model architectures and game environments to understand how self-play develops transferable reasoning capabilities. We train on three games from TextArena (Guertler et al., 2025): TicTacToe (spatial reasoning), Kuhn Poker (probabilistic reasoning), and Simple Negotiation (strategic optimization). Models include Qwen3-4B/8B-Base (Yang et al., 2025), Llama-3.1-8B-Instruct (Dubey et al., 2024), and Octothinker-8B-Base (Wang et al., 2025a). Training spans 400 steps with 128 samples per step on 8 H100 GPUs, using Adam optimizer with learning rate $1\times 10^{-6}$ and temperature 1.0. We evaluate on eight reasoning benchmarks (MATH500, OlympiadBench, Minerva Math, AIME24/25, AMC23, GPQA-Diamond, MMLU-Pro) and seven out-of-distribution games. Complete implementation details are in Appendix D.
Self-play on games transfers to improve reasoning. The central results of this paper are shown in Table 1, which demonstrates that multi-game SPIRAL training achieves up to 10.5% improvement on reasoning benchmarks without domain-specific data. with Qwen3-4B-Base improving from 34.0% to 44.5% average performance (+10.5+10.5%). We also compare with supervised fine-tuning (SFT) on 25,000 expert game trajectories, generated by Qwen3-32B models, which improves performance on several benchmarks including AIME24 and AIME25, revealing that games themselves contain skills relevant to reasoning. However, SPIRAL consistently outperforms SFT across all 8 benchmarks, demonstrating that self-play discovers more effective reasoning strategies than imitating expert demonstrations.
4.1 Understanding Why SPIRAL Works
Figure 4: Evolution of reasoning patterns during SPIRAL training and their transfer to mathematical reasoning. We track three core reasoning patterns (Pattern Recognition, Expected Value Calculation, and Case-by-Case Analysis) across 290 game trajectories and 46,792 math solutions. Left: Patterns increase via training on games, with Expected Value Calculation reaching 78% by late training. Middle: These patterns transfer to mathematical reasoning with varying effectiveness: Case-by-Case Analysis maintains high transfer (72% to 71%), Pattern Recognition shows amplification (35% to 45%), while Expected Value Calculation transfers more selectively (78% to 28%). Right: Math benchmark scores improve from 31.2 to 39.6 as these reasoning patterns develop, demonstrating that game-learned strategies enhance mathematical problem-solving capabilities.
Emergent reasoning patterns learned on games transfer to math questions. To understand how games can improve reasoning performance, we analyzed chain-of-thought patterns using GPT-4.1 (OpenAI, 2025) to classify 290 game trajectories and 46,792 math solutions. Table 2 illustrates three core patterns that emerge during gameplay and transfer to mathematics. Figure 4 tracks their evolution. In the late training stage, the transfer from games to math is clear: Case-by-Case Analysis transfers near-perfectly (72% to 71%), Pattern Recognition amplifies (35% to 45%), while Expected Value Calculation transfers selectively (78% to 28%). The success of this transfer may stem from two factors. This transfer likely works because evolving opponents prevent memorization while games teach pure reasoning strategies without requiring mathematical knowledge.
@@ 4.1 Understanding Why SPIRAL Works
Table 2: Representative reasoning patterns and their manifestations in games and mathematics. We use an LLM-as-a-judge framework to evaluate the frequency of these patterns over training.
@@
Case-by-Case Analysis <output>
@@
Since losing 1 chip (fold) is better than losing 2 chips (call and lose), the best action is to $\boxed{fold}$.
</output>
<output> To solve $m^{2}+2\cdot 3^{n}=m(2^{n+1}-1)$:
@@
Case $n=4$: $m^{2}+162=31m$, no integer solutions
Case $n=5$: $m^{2}+486=63m$ gives $m=\boxed{9,54}$ (valid) </output>
Expected Value Calculation <output> With card J against opponent’s bet:
@@
Since EV(fold) >> EV(call), $\boxed{fold}$ is optimal. </output>
<output> For average of $A+BC+DEF$ over permutations:
@@
Total = $\frac{1}{3}+\frac{1}{9}-\frac{1}{27}=\boxed{-\frac{7}{27}}$ </output>
Pattern Recognition <output> Previous rounds: Player 0 had K both times and won. Pattern identified: Player 0 likely has strong cards or bluffs consistently. Given I have J (weak), betting might exploit their aggressive calling pattern. </output> <output> Sum of three consecutive integers (n1)+n+(n+1)=3n(n-1)+n+(n+1)=3n.
@@
Smallest: $k=\boxed{3}\Rightarrow 3n=27$ </output>
Adaptive curriculum beats static opponents. Self-play creates an automatic curriculum that adapts to model capabilities. Figure 5 compares self-play against fixed opponents (Random, Mistral-Small-3, Gemini-2.0-Flash-Lite [^1] Accessed via https://openrouter.ai/google/gemini-2.0-flash-lite-001.).
@@
Random opponents cause collapse: although they provide randomized rewards with positive expected value, similar to spurious rewards that might upweight certain base model behaviors (Shao et al., 2025), we observe this is insufficient to improve performance. Fixed model opponents enable initial learning but plateau once exploitable strategies are found. That fixed opponents like Gemini yield smaller gains reveals the effects are not merely from learning game mechanics or spurious rewards (Shao et al., 2025), but specifically from the adaptive curriculum. Unlike static baselines, self-play’s continuously evolving challenge forces genuine reasoning development rather than exploitation of static patterns. Table 3 confirms this: self-play maintains 50-52% win rates while fixed-opponent training rises from 0% to 62.5%, indicating exploitation rather than continued learning.
Figure 5: Performance comparison of self-play training and fixed-opponent baselines. All evaluations are averaged over multiple games/benchmarks (see Sec.˜D.3). Mistral Opponent refers to against Mistral-Small-3; Gemini Opponent refers to against Gemini-2.0-Flash-Lite.
Different games develop complementary skills. Each game cultivates distinct cognitive abilities that transfer to related domains; specifically, we selected TicTacToe to target spatial reasoning, Kuhn Poker for probabilistic inference, and Simple Negotiation for strategic optimization. Table 4 tests how well agents trained on a specific game (‘specialists’) transfer to novel out-of-distribution (OOD) games. We find specialists transfer effectively to similar out-of-distribution games: TicTacToe specialists achieve 56.0% on Snake (spatial), Poker specialists dominate Pig Dice at 91.7% (probabilistic), Negotiation specialists win 55.8% on Truth and Deception (strategic). Multi-game training combines these skills synergistically, as shown in Table 5, which shows win-rate against Gemini-2.0-Flash [^2] Accessed via https://openrouter.ai/google/gemini-2.0-flash-001.. Multi-game agents achieve 59.5% average performance, outperforming all single-game specialists (best: 52.9%), using a fixed budget of 400 training steps determined by convergence analysis.
@@
Table 5: Multi-game training achieves competitive performance across all training games while excelling at novel composite challenges. All win rates shown are against Gemini-2.0-Flash as a fixed opponent. The multi-game model outperforms all specialists on average, demonstrating that diverse game training develops more flexible reasoning.
@@
Model TicTacToe KuhnPoker Simple Negotiation Snake Pig Dice Truth and Deception Average
Base Model 17.5% 21.5% 15.6% 7.8% 0.2% 49.6% 18.7%
Random Policy 24.5% 31.3% 8.2% N/A N/A N/A N/A
Instruct Model 52.7% 48.5% 46.2% 25.2% 97.6% 75.5% 57.6%